Problem 13-12 Calculating Portfolio Betas [LO4]

You own a portfolio equally invested in a risk-free asset and two stocks. If one of the stocks has a beta of 1.11 and the total portfolio is equally as risky as the market, what must the beta be for the other stock in your portfolio? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Market's Beta = 1

Portfolio Beta = [weight(i) * beta(i)]

1 = [0.5 * 1.11] + [0.5 * beta(other stock)]

1 = 0.555 + [0.5 * beta(other stock)]

0.5 * beta(other stock) = 1 - 0.555

beta(other stock) = 0.445 / 0.5 = 0.89